OpenStudy (anonymous):

What is the input value for the following function if the output value is 1.2? y = x - 0.4 0.8 8.4 1.6 4.2

Parth (parthkohli):

x is the input.

OpenStudy (blues):

Here the output value they give you is the value for y. So you substitute that in for y and solve for x, the input...

Parth (parthkohli):

1.2 = x - 0.4 x = ?

OpenStudy (anonymous):

or 1.2+0.4= ?

Parth (parthkohli):

Hmm...you need not answer please